OpenStudy (anonymous):

Does anyone know how to do this? "The current (I) in an electrical conductor varies inversely as the resistance (r) of the conductor. The current is 2 amperes when the resistance is 960 ohms. What is the current when the resistance is 480 ohms?"

OpenStudy (anonymous):

do you know the letter for current? the letter for voltage is V, for resistance it R... and current is?

OpenStudy (anonymous):

\[I \alpha \frac{ 1 }{ R } \] this just shows the inverse proportionality of I and R. This is expressed mathematically as \[I= \frac{ k }{ R }\]

OpenStudy (anonymous):

I just copied and pasted the question.

OpenStudy (anonymous):

they use I for current

OpenStudy (anonymous):

\[I=\frac{ V }{ R }\] in the first case we have given I=2 A, R=960 Ohms

OpenStudy (anonymous):

So, you need to find k first. To do this you were given that I= 2 amperes when R = 960 ohms. Plugging in these values in the equation above, \[k= (2 amperes)* (960 ohms)\], so k= 1920 amperes *ohms.

OpenStudy (anonymous):

Now to get I when R is 480 you just have to plug in R= 480 ihms and k= 1920 ohms*amperes in the formula. So I= (1920 ohms*amperes)/ 480 ohms. So I= 4 amperes.

OpenStudy (anonymous):

I appreciate the help. Thank you

OpenStudy (anonymous):

You're welcome!